# ELECTRA and GPT-4o: Cost-Effective Partners for Sentiment Analysis

### v1

@@ 6 Results
B3 4o-mini Baseline (Prompt only, Zero shot) 79.52 80.34 81.12 81.00 77.35 77.92 70.67 80.05
B4 4o Baseline (Prompt only, Zero shot) 80.14 80.74 81.12 80.94 80.22 80.56 72.20 80.45
@@
E7 4o-mini Large FT Prompt, Probabilities 83.01 83.60 86.44 86.36 78.92 79.17 71.53 80.54 53.26 0.64
E8 4o-mini Large FT Prompt, Label, Probabilities 83.43 84.12 87.02 86.94 79.72 80.14 71.03 80.81 53.26 0.64
@@
E11 4o-mini FT-M Minimal fine-tune 86.55 87.00 89.70 89.58 86.97 87.08 75.62 82.76 16.60 0.19
E12 4o-mini FT Base FT Prompt, Label, FT w/prompt 81.42 81.90 84.77 84.78 74.49 74.72 70.95 79.55 42.88 0.53
E13 4o-mini FT-L Base FT Prompt, Label, FT w/prompt, label 82.02 82.53 85.24 85.11 78.15 78.47 71.68 79.64 49.31 0.60
E14 4o Large FT Prompt, Label 81.57 81.98 84.03 83.86 77.69 77.92 72.94 80.23 53.26 0.65
E15 4o Large FT Prompt, Label, Examples (Few shot) 83.09 83.66 86.01 85.86 81.53 81.81 72.06 80.68 53.26 0.64
E16 4o Large FT Prompt, Label, Balanced Ex. (Few shot) 82.99 83.55 85.87 85.69 80.89 81.11 72.15 80.86 53.26 0.64
E17 4o Large FT Prompt, Probabilities 82.65 83.25 86.05 85.97 77.67 77.92 71.16 80.54 53.26 0.64
E18 4o Large FT Prompt, Label, Probabilities 83.08 83.71 86.44 86.33 79.23 79.58 71.48 80.77 53.26 0.64
E19 4o Large FT Prompt, Label, Probs, Examples 82.74 83.35 86.32 86.22 77.76 78.06 70.98 80.41 53.26 0.64
E20 4o FT Fine-tune w/prompt 86.83 87.43 90.44 90.36 88.34 88.47 73.08 82.31 276.24 3.18
E21 4o FT-M Minimal fine-tune 86.99 87.57 90.57 90.50 89.00 89.17 73.99 82.26 138.37 1.59
E22 4o FT Large FT Prompt, Label, FT w/prompt 83.82 84.47 87.80 87.72 79.49 79.86 71.18 80.68 329.50 3.93
E23 4o FT-L Large FT Prompt, Label, FT w/prompt, label 84.23 84.82 87.74 87.64 80.55 80.83 72.63 81.54 383.10 4.55
@@
  • 1 GPT fine-tuning types: FT = fine-tune all layers with prompt, FT-M = minimal format without prompt, FT-L = with prompt including ELECTRA label
  • 2 Merged dataset: Combination of test splits from DynaSent R1/R2 and SST-3
  • 3 Cost: FT = Fine-tuning cost, no inference-time API charges. Ratio is FT cost divided by F1 score.
Our experiments revealed significant differences in performance across baseline, fine-tuning, and collaborative scenarios. See Table 6 for the results.
Baselines. Regarding baselines, both GPT models outperformed the ELECTRA classifiers, with GPT-4o achieving 80.14 macro F1 and GPT-4o-mini scoring 79.52, compared to ELECTRA Base (69.65) and Large (67.88). This demonstrates the strong zero-shot capabilities of the GPT models.
Fine-tuning. Fine-tuning improved performance across all models. ELECTRA Base’s macro F1 increased from 69.65 to 79.29, while ELECTRA Large showed even greater gains, improving from 67.88 to 82.36. This improvement is the result of fine-tuning all layers—the baselines had the same classifier head. The fine-tuned GPT models had the highest scores (see Figure 2), with GPT-4o-mini FT rising from 79.52 to 86.77, and GPT-4o FT-M achieving 86.99 with the minimal template.
Sharing Predictions. The effect of adding ELECTRA predictions to GPT prompts depended on if the GPT model was fine-tuned (see Figure 1). Sharing ELECTRA Base predictions with GPT-4o-mini (not fine-tuned) significantly improved the macro F1 from 79.52 to 82.74 (p < 0.0001, McNemar’s test), a +3.22 gain. There was an even greater gain of +3.97 points when ELECTRA Large predictions were shared (from 79.52 to 83.49, p < 0.0001). Similarly, including ELECTRA Large predictions with GPT-4o improved the macro F1 from 80.14 to 81.57 (p = 0.0106), a smaller +1.43 gain.
However, sharing ELECTRA predictions with fine-tuned GPT models actually decreased performance. GPT-4o-mini’s macro F1 dropped from 86.77 to 81.42 with ELECTRA Base predictions, and to 82.02 when fine-tuned with the predictions included in the prompt. Similarly, GPT-4o fell from 86.83 to 83.82 with ELECTRA Large predictions, and to 84.23 when fine-tuned with them.
Few-shot Examples. Some contexts performed better than others for specific model combinations (see Figure 3). For GPT-4o with ELECTRA Large FT, providing few-shot examples in addition to the predicted label significantly increased the macro F1 on the Merged test set from 81.57 to 83.09 (p = 0.0001, McNemar’s test), a +1.51 gain. The difference was even greater for DynaSent R2, increasing from 77.69 to 81.53, +3.84 gain.
Figure 1: Change in F1 from Adding Predictions
Figure 2: Macro F1 Scores on Merged Dataset by Experiment
There was very little difference between showing the top five similar examples, compared with the top six balanced examples (ensuring two examples for each class). However, the top five approach had slightly higher scores, except for GPT-4o-mini on DynaSent R2, and GPT-4o on SST-3—both of which represent more challenging sentiment.
Figure 3: Impact of Context on F1 Score by Model
Sharing Probabilities. For GPT-4o with ELECTRA Large, using probabilities instead of the predicted label resulted in a slight increase on the Merged test macro F1 (from 81.57 to 82.65, p = 0.0049), a +1.08 gain. Providing both the label and probabilities was even more effective, boosting the score from 81.57 to 83.08 (p = 0.0001), a +1.51 gain. The difference was even greater for DynaSent R1, increasing from 84.03 to 86.44, a +2.41 gain.
For GPT-4o-mini with ELECTRA Large FT on the Merged test, there was no context better than just the predicted label (see Figure 3). However, adding probabilities in addition to the label increased the macro F1 on DynaSent R2 from 76.19 to 79.72, a +3.53 gain.
Datasets. Performance also varied across datasets. GPT-4o FT-M achieved the top scores on DynaSent R1 (90.57 macro F1) and DynaSent R2 (89.00). Surprisingly, GPT-4o-mini FT—the smaller model—performed best on SST-3 with a 75.68 macro F1, exceeding even GPT-4o FT’s performance of 73.99.
Figure 4: Incorrect Predictions by Label and Experiment
7 Analysis
However, following ELECTRA’s predictions had mixed results. When GPT-4o-mini changed its decision and followed ELECTRA Base FT, it was correct 548 times and wrong 412 times (+136 net improvement, 57.08% success rate). When GPT-4o changed its decision and followed ELECTRA Large FT, it was correct 521 times and wrong 481 times (+40 net improvement, 52% success rate).
H2. Improvement would be less for fine-tuned GPTs. H2 was supported more strongly than anticipated. For a fine-tuned GPT model, including the ELECTRA prediction actually decreased performance (see Figure 1). Initially, we thought this was because the fine-tuning context did not include the ELECTRA prediction in the prompt. But we still saw a decrease in performance (although less) when it was included. This suggests the fine-tuning process for GPT-4o/4o-mini is superior when it’s simpler—focused on pure input to output.
H3. Format of prediction would impact performance. The significant improvement in GPT-4o’s performance when ELECTRA Large predictions were replaced with probabilities as percentages supports H3. For GPT-4o-mini, probabilities decreased performance on the Merged dataset and DynaSent R1, but improved performance for DynaSent R2 and SST-3—the more challenging sentiment examples. In almost all cases, it was better to include both the label and probabilities together.
H4. Including examples would improve performance. The significant improvement in GPT-4o’s performance when few-shot examples were included with the predicted label strongly supports H4. Of all the contexts explored with GPT-4o (not fine-tuned), including the top five examples yielded the best score for DynaSent R2 (81.53 macro F1). Similarly, with GPT-4o-mini (not fine-tuned), the top examples produced its best performance with SST-3 (71.98 macro F1). It appears the examples were able to help the models make better decisions with the more challenging datasets.
Overall, adding probabilities, or adding examples, improved performance over the predicted label alone for GPT-4o. However, including all three (label, probabilities, examples) was not as good as either pairing (although still better than the label alone). Perhaps the large amount of tokens impacted the signal to noise ratio.
8 Conclusion
This research investigated collaborative approaches to sentiment classification between bidirectional transformers and LLMs. Our results show that augmenting prompts with predictions from a fine-tuned ELECTRA can significantly improve performance when the GPT model is not fine-tuned—up to +3.97 points of gain in the macro F1 score. Including probabilities or similar examples enhanced performance for GPT-4o, especially on challenging datasets. However, this collaborative benefit disappeared when the GPT models were fine-tuned.
Our findings offer several cost-effective paths for sentiment analysis projects. For organizations that can fine-tune via API, GPT-4o-mini FT offers nearly equivalent performance to GPT-4o FT-M (86.77 vs 86.99 macro F1) at 76% lower cost (\$0.38 vs \$1.59/F1 point). For those with data privacy concerns or resource constraints, GPT-4o-mini with ELECTRA Base FT had the best cost/performance ratio (\$0.12/F1 point). Projects that need to stay completely local can fine-tune ELECTRA Large, which outperformed both base GPT models.
9 Limitations
The cost/performance evaluation only considered the fine-tuning costs to achieve the reported macro F1 on the test set. In practice, there would be ongoing costs for inference time API calls.
Appendix A ELECTRA Fine-tuning Details
@@
Appendix B OpenAI Fine-tuning Templates
@@
Appendix C DSPy Prompt Signature Examples

### v2

@@ Appendix A Results of Individual Runs
B3 4o-mini Baseline (Zero shot) 79.52 80.34 81.12 81.00 77.35 77.92 70.67 80.05
B4 4o Baseline (Zero shot) 80.14 80.74 81.12 80.94 80.22 80.56 72.20 80.45
@@
E7 4o-mini Large FT Prompt, Probs 83.01 83.60 86.44 86.36 78.92 79.17 71.53 80.54 53.26 0.64
E8 4o-mini Large FT Prompt, Label, Probs 83.43 84.12 87.02 86.94 79.72 80.14 71.03 80.81 53.26 0.64
@@
E11 4o-mini FT 5 Fine-tune w/prompt (5 epochs) 84.95 85.27 87.83 87.67 85.93 85.97 75.60 81.13 165.75 1.95
E12 4o-mini FT-M Minimal fine-tune 86.55 87.00 89.70 89.58 86.97 87.08 75.62 82.76 16.60 0.19
E13 4o-mini FT Base FT Prompt, Label, FT w/prompt 81.42 81.90 84.77 84.78 74.49 74.72 70.95 79.55 42.88 0.53
E14 4o-mini FT-L Base FT Prompt, Label, FT w/prompt, label 82.02 82.53 85.24 85.11 78.15 78.47 71.68 79.64 49.31 0.60
E15 4o-mini FT-L 5 Base FT Prompt, Label, FT w/prompt, label (5 epochs) 83.88 84.27 86.74 86.61 81.62 81.81 75.00 81.27 207.64 2.48
E16 4o-mini FT Large FT Fine-tune w/prompt 84.00 84.58 87.65 87.58 80.37 80.69 72.46 80.95 86.41 1.03
E17 4o-mini FT-L Large FT Fine-tune w/prompt, label 84.16 84.70 87.65 87.56 80.89 81.11 73.29 81.22 92.84 1.10
E18 4o-mini FT-L 5 Large FT Fine-tune w/prompt, label (5 epochs) 84.87 85.25 87.87 87.75 83.80 83.89 75.58 81.63 251.17 2.96
E19 4o Large FT Prompt, Label 83.18 83.68 85.71 85.56 81.98 82.22 73.44 81.09 53.26 0.64
E20 4o Large FT Prompt, Label, Examples (Few shot) 83.09 83.66 86.01 85.86 81.53 81.81 72.06 80.68 53.26 0.64
E21 4o Large FT Prompt, Label, Balanced Ex. (Few shot) 82.99 83.55 85.87 85.69 80.89 81.11 72.15 80.86 53.26 0.64
E22 4o Large FT Prompt, Probs 82.65 83.25 86.05 85.97 77.67 77.92 71.16 80.54 53.26 0.64
E23 4o Large FT Prompt, Label, Probs 83.08 83.71 86.44 86.33 79.23 79.58 71.48 80.77 53.26 0.64
E24 4o Large FT Prompt, Label, Probs, Examples 82.74 83.35 86.32 86.22 77.76 78.06 70.98 80.41 53.26 0.64
E25 4o FT Fine-tune w/prompt 86.83 87.43 90.44 90.36 88.34 88.47 73.08 82.31 276.24 3.18
E26 4o FT-M Minimal fine-tune 86.99 87.57 90.57 90.50 89.00 89.17 73.99 82.26 138.37 1.59
E27 4o FT Large FT Fine-tune w/prompt 83.82 84.47 87.80 87.72 79.49 79.86 71.18 80.68 329.50 3.93
E28 4o FT-L Large FT Fine-tune w/prompt, label 84.23 84.82 87.74 87.64 80.55 80.83 72.63 81.54 383.10 4.55
Table 8: Round 2 Results
Merged DynaSent R1 DynaSent R2 SST-3 Cost (\$)
ID GPT1 ELECTRA Description F1 Acc F1 Acc F1 Acc F1 Acc FT /F1
B1 Base Baseline, Classifier head 69.37 69.57 70.75 71.11 61.19 61.25 60.34 69.77 0.65 0.01
B2 Large Baseline, Classifier head 67.99 68.15 69.67 70.03 59.78 59.72 58.21 67.83 2.51 0.04
B3 4o-mini Baseline (Zero shot) 79.29 80.15 81.19 81.08 76.69 77.36 69.30 79.55
B4 4o Baseline (Zero shot) 79.80 80.47 80.77 80.61 80.05 80.42 71.96 80.27
E1 Base FT Fine-tune all layers 78.98 79.46 82.13 82.19 69.51 69.72 68.13 78.19 9.73 0.12
E2 Large FT Fine-tune all layers 83.16 83.71 86.53 86.44 78.36 78.61 72.63 80.91 53.26 0.64
E3 4o-mini Base FT Prompt, Label 82.26 82.92 86.29 86.25 74.47 74.86 70.03 80.14 9.73 0.12
E4 4o-mini Large FT Prompt, Label 84.10 84.73 87.90 87.83 79.52 79.86 72.54 81.27 53.26 0.63
E5 4o-mini Large FT Prompt, Label, Examples (Few shot) 83.63 84.18 87.13 87.03 80.29 80.56 72.68 80.72 53.26 0.64
E6 4o-mini Large FT Prompt, Label, Balanced Ex. (Few shot) 83.28 83.81 86.71 86.58 80.12 80.42 72.51 80.41 53.26 0.64
E7 4o-mini Large FT Prompt, Probs 83.53 84.07 86.76 86.67 79.89 80.14 72.98 81.13 53.26 0.64
E8 4o-mini Large FT Prompt, Label, Probs 83.88 84.47 87.41 87.33 80.23 80.56 72.53 81.09 53.26 0.64
E9 4o-mini Large FT Prompt, Label, Probs, Examples 83.46 84.03 87.00 86.89 79.28 79.58 72.39 80.81 53.26 0.64
E10 4o-mini FT Fine-tune w/prompt 86.62 87.09 89.44 89.33 87.08 87.22 75.98 83.39 33.15 0.38
E11 4o-mini FT 5 Fine-tune w/prompt (5 epochs) 84.76 85.07 87.64 87.47 86.50 86.53 75.15 80.68 165.75 1.96
E12 4o-mini FT-M Minimal fine-tune 86.47 86.89 89.44 89.31 87.28 87.36 75.86 82.81 16.60 0.19
E13 4o-mini FT Base FT Prompt, Label, FT w/prompt 80.69 81.24 84.57 84.58 71.62 71.94 68.45 78.82 42.88 0.53
E14 4o-mini FT-L Base FT Prompt, Label, FT w/prompt, label 81.65 82.25 85.15 85.03 76.43 76.81 69.71 79.50 49.31 0.60
E15 4o-mini FT-L 5 Base FT Prompt, Label, FT w/prompt, label (5 epochs) 83.46 83.89 86.02 85.89 80.75 80.97 75.04 81.58 207.64 2.49
E16 4o-mini FT Large FT Fine-tune w/prompt 83.87 84.44 87.49 87.42 79.97 80.28 72.46 80.95 86.41 1.03
E17 4o-mini FT-L Large FT Fine-tune w/prompt, label 84.07 84.59 87.51 87.42 80.61 80.83 73.38 81.22 92.84 1.10
E18 4o-mini FT-L 5 Large FT Fine-tune w/prompt, label (5 epochs) 84.79 85.15 87.63 87.50 84.94 85.00 75.59 81.36 251.17 2.96
E19 4o Large FT Prompt, Label 83.20 83.69 85.71 85.56 82.13 82.36 73.52 81.09 53.26 0.64
E20 4o Large FT Prompt, Label, Examples (Few shot) 83.48 84.00 86.21 86.06 81.43 81.67 73.85 81.40 53.26 0.64
E21 4o Large FT Prompt, Label, Balanced Ex. (Few shot) 83.38 83.89 86.14 85.97 81.18 81.39 73.61 81.31 53.26 0.64
E22 4o Large FT Prompt, Probs 83.32 83.87 86.68 86.58 79.16 79.44 72.63 80.90 53.26 0.64
E23 4o Large FT Prompt, Label, Probs 83.54 84.10 86.93 86.81 79.69 80.00 72.85 81.04 53.26 0.64
E24 4o Large FT Prompt, Label, Probs, Examples 83.34 83.89 86.73 86.61 79.18 79.44 72.68 80.90 53.26 0.64
E25 4o FT Fine-tune w/prompt 86.74 87.32 90.48 90.42 87.94 88.06 73.09 82.04 276.24 3.18
E26 4o FT-M Minimal fine-tune 86.99 87.57 90.57 90.50 89.00 89.17 73.99 82.26 138.37 1.59
E27 4o FT Large FT Fine-tune w/prompt 84.24 84.84 87.99 87.89 80.52 80.83 72.81 81.18 329.50 3.91
E28 4o FT-L Large FT Fine-tune w/prompt, label 84.50 85.04 87.87 87.75 82.01 82.22 73.56 81.54 383.10 4.53
@@
  • 1 GPT fine-tuning types: FT = fine-tune all layers with prompt, FT-M = minimal fine-tune format without prompt, FT-L = fine-tune with prompt including ELECTRA label, FT 5 = fine-tune for 5 epochs
Appendix B ELECTRA Fine-tuning Details
@@
Appendix C GPT Fine-tuning Details
Table 10: GPT Fine-Tune Configuration
Setting 4o-mini
FT-M
4o-mini
FT
4o-mini
FT-L Base 1
4o-mini
FT-L Base 5
4o-mini
FT-L Large 1
4o-mini
FT-L Large 5
4o
FT-M
4o
FT
4o
FT-L Large 1
GPT Model 4o-mini 4o-mini 4o-mini 4o-mini 4o-mini 4o-mini 4o 4o 4o
ELECTRA Model None None Base FT Base FT Base FT Base FT None None Large FT
Code FT-M FT FT-L FT-L FT-L FT-L FT-M FT FT-L
Code Meaning Minimal
Format
Fine-Tune Fine-Tune w/Label Fine-Tune w/Label Fine-Tune w/Label Fine-Tune w/Label Minimal
Format
Fine-Tune Fine-Tune w/Label
Format JSON
(No Prompt)
DSPy Prompt DSPy Prompt + Label DSPy Prompt + Label DSPy Prompt + Label DSPy Prompt + Label JSON
(No Prompt)
DSPy Prompt DSPy Prompt + Label
Source OpenAI OpenAI OpenAI OpenAI OpenAI OpenAI OpenAI OpenAI OpenAI
Source Model ID gpt-4o-mini-2024-07-18 gpt-4o-mini-2024-07-18 gpt-4o-mini-2024-07-18 gpt-4o-mini-2024-07-18 gpt-4o-mini-2024-07-18 gpt-4o-mini-2024-07-18 gpt-4o-2024-08-06 gpt-4o-2024-08-06 gpt-4o-2024-08-06
Dataset Merged Merged Merged Merged Merged Merged Merged Merged Merged
Train Size 102,097 102,097 102,097 102,097 102,097 102,097 102,097 102,097 102,097
Validation Size 5,421 5,421 5,421 5,421 5,421 5,421 5,421 5,421 5,421
Fine-Tuning Date 2024-10-23
Total Job Time 01:17:07 01:07:05 01:09:00 01:09:00 01:09:00 01:09:00 01:31:06 01:43:43 01:47:49
LR Multiplier 1.8 1.8 1.8 1.8 1.8 1.8 2.0 2.0 2.0
Seed 42 42 42 42 42 42 42 42 42
Batch Size 68 68 68 68 68 68 68 68 68
Epochs 1 1 1 1 1 1 1 1 1
Train Loss 0.1250 0.1383 0.1286 0.1286 0.1286 0.1286 0.1185 0.1150 0.1075
Validation Loss 0.1112 0.1114 0.1429 0.1429 0.1429 0.1429 0.1054 0.1056 0.1242
Trained Tokens 5,534,849 11,049,720 13,193,757 13,193,757 13,193,757 13,193,757 5,534,849 11,049,720 13,193,757
Rate \$3/1M tokens \$3/1M tokens \$3/1M tokens \$3/1M tokens \$3/1M tokens \$3/1M tokens \$25/1M tokens \$25/1M tokens \$25/1M tokens
Cost \$16.60 \$33.15 \$39.58 \$39.58 \$39.58 \$39.58 \$138.37 \$276.24 \$329.84
Appendix D OpenAI Fine-tuning Templates
@@
Appendix E DSPy Prompt Signature Examples